OpenStudy (anonymous):

Hi all. I am dealing with e problem and I hope you can help me. I have already proved this: Let us suppose that integers m and n can be written as sum of squares of two integers. Prove that m*n can also be written as sum of squares of two integers. Now I am trying to prove that m*n can also be written as sum of squares of four integers. Can you help me on this? Thank You for reading my post.

OpenStudy (amistre64):

let m=a^2+b^2 let n=x^2+y^2 mn = (a^2+b^2)(x^2+y^2)

OpenStudy (amistre64):

mn = (ax)^2 + (ay)^2 + (bx)^2 + (by)^2

OpenStudy (amistre64):

how did you prove the first one?

OpenStudy (anonymous):

a^2+b^2=m c^2+d^2=n (ac)^2+(ad)^2+(bc)^2+(bd)^2=mn (ac)^2+2acbd+(bd)^2-2acbd+(ad)^2-2adbc+(bc)^2+2adbc=mn (ac+bd)^2+(ad-bc)^2=mn That's how i proved it:)

OpenStudy (amistre64):

cool, so in your proofing of the first one, youve actually proven the second one

OpenStudy (anonymous):

But the problem is that m should be equal to a^2+b^2+c^2+d^2 and n should we equal to f^2+l^2+g^2+s^2.

OpenStudy (amistre64):

that is not what is stated originally Suppose: m and n can be written as sum of squares of two integers m = (a^2 + b^2) is the sum of two integers square n = (x^2 + y^2) is the sum of two integers square Prove that: m*n can be written as sum of squares of four integers. if we let m*n = k, such that \(k = s^2+t^2+u^2+v^2\)

OpenStudy (amistre64):

the inital conditions for m and n havent changed as far as I am aware

OpenStudy (anonymous):

Oh it's my mistake,I thought i have mentioned that :S

OpenStudy (amistre64):

so to be clear: are we saying that the inital conditions for m and n have changed? or have they remained the same and youve simply misread the problem?

OpenStudy (anonymous):

Hi all. I am dealing with a problem and I hope you can help me. I have already proved this: Let us suppose that integers m and n can be written as sum of squares of two integers. Prove that m*n can also be written as sum of squares of two integers. Now I am trying to prove this:Let us suppose that integers m and n can be written as sum of squares of four integers. Prove that m*n can also be written as sum of squares of four integers.

OpenStudy (amistre64):

a+b+c+d w+x+y+z ---------- aw+bw+cw+dw ax+bx+cx+dx ay+by+cy+dy az+bz+cz+dz looks to me like you just have to work out 4 completed squares

OpenStudy (amistre64):

a+b x+y ----- ax+bx+2abxy ay+by -2axby a+b+c+d w+x+y+z ---------- aw+bw +2abwx +cw+dw +2cdwx ax+bx -2abwx +cx+dx -2cdwx ay+by +2abyz +cy+dy +2cdyz az+bz -2abyz +cz+dz -2cdyz

OpenStudy (amistre64):

i spose that 8 integers :)

OpenStudy (amistre64):

at least thats better than 16 integers that we started with rename the 8 integers in terms that can be reduced to 4 terms is an idea